if y=7.5 when x=0.5, find y when x= -0.3
Answer by stanbon(75887)

if y=7.5 when x=0.5, find y when x= -0.3
---
Use a proportion:
y/x = y/x
---------
y/-0.3 = 7.5/0.5
-----
y = -0.3*15
---
y = -0.45
================
Cheers,
Stan H.
================